OpenStudy (girlgoyleh.):

Which answer shows this equation in standard form? 8 – 4(x – y) = –6x + 6 A. 2x – 4y = –2 B. –10x + 4y = –2 C. 2x + 4y = 2 D. 2x + 4y = –2

OpenStudy (zale101):

8 – 4(x – y) = –6x + 6 Distribute and combine like terms. You know where to distribute?

OpenStudy (girlgoyleh.):

@Zale101

OpenStudy (girlgoyleh.):

no i dont

OpenStudy (zale101):

a(b+c)=ab+ac